OpenStudy (anonymous):

x+5=-6x-16 Please help I dont get these. I have 5 total What is the value of x?

OpenStudy (only_hanaaa):

well first you need to isolate the x

OpenStudy (anonymous):

okay but how do i do that with the =

OpenStudy (pawanyadav):

x-(-6x)=-16-5 7x=-21 x=-3

OpenStudy (ivaltino):

7x+5=−16

OpenStudy (anonymous):

Thanks!

OpenStudy (pawanyadav):

Welcome

OpenStudy (ivaltino):

7x+5=−16 7x=−5−16 7x=−21 x=−3

OpenStudy (anonymous):

You have to have x on the same side , so it wont be on two sides. if that helps at all.

OpenStudy (pawanyadav):

x-(-6x)=-16-5 7x=-21 x=-3